on two occasions last month,Sam rented a car from the same place and same model.He paid 210 and drove 625 miles the first trip.Second trip he paid 133.50 and drove 370 miles.
After plotting these two points on a graph and drawing the line that goes through
them, answer the following:
What is the equation of the line in the slope intercept form of the equation?
In your own words, explain what question the slope of the line answers for Sam.
What does the y intercept tell Sam and how does it help him?
Write a one paragraph memo explaining to Sam the answers to his question and
how knowledge of algebra and graphing helped you to find the answers.

Answer by stanbon(75887)

on two occasions last month,Sam rented a car from the same place and same model.He paid 210 and drove 625 miles the first trip.Second trip he paid 133.50 and drove 370 miles.
After plotting these two points on a graph and drawing the line that goes through them, answer the following:
---
Using (625 mi ; 210 dollars) and (370 miles ; 133.50 dollars)
---
Slope = (210-133.5)/(625-370) = 0.3
----
Form: y = mx + b
210 = 0.3*625+b
b = 22.5
---
Cost = 0.3(miles) + 22.5
---------
What is the equation of the line in the slope intercept form of the equation?
------
In your own words, explain what question the slope of the line answers for Sam.
For each mile he drives the cost rises 30 cents.
-----------------------------
What does the y intercept tell Sam and how does it help him?
The initial cost of renting the car was $22.50
===================
I'll leave the discussion paragraph to you.
=======================
Cheers,
Stan H.
================